MadisonMarvel:

Which expression is equivalent to 144^3/2

supie:

Choices?

MadisonMarvel:

A.216 B.1728 C. 3 √(12) D.3 √(72)

jhonyy9:

so this 2 on denominator of exponent mean that this is square root of 144^3 \[\sqrt{(144)^3} = \sqrt{(12^2)^3} = [(12)^2]^{(3/2)} = ?\] do you can continue it ?